OpenStudy (anonymous):

need this ASAP: a rectangular enclosure must have an area of at least 800yd^2. If 180yd of fencing is to be used, and the width cannot exceed the length, withing what limits must the width of the enclosure be?

OpenStudy (anonymous):

wait so what am i doing? sorry i dont have much time

OpenStudy (xishem):

Since the area is 800yd^2 or greater...\[(l*w)>=A_{\min}\]And...\[2l+2w=length_{fence}\]...\[2l+2w=180 \rightarrow l=\frac{180-2w}{2} \rightarrow l=90-w\]Substitute back into the area equation, including the explicit value for the minimum area...\[l*w >= A_{\min} \rightarrow (90-w)*w>=800\]\[90w-w^2 >= 800 \rightarrow w^2-90w+800 >= 0\]I'm not going to take you any further than that. You'll have to work it out on your own. Try it and see how far you can get.

OpenStudy (xishem):

Sorry, it should be...\[w^2-90w+800 <= 0\]

OpenStudy (anonymous):

10<=x<=80.......?

OpenStudy (anonymous):

@Xishem ?

jimthompson5910 (jim_thompson5910):

you're close, but keep in mind that "the width cannot exceed the length" So \(\Large W \le L\)